15 January 2016 UT: 1st half D night

Manager: J. Little
OSA: G. Bechetti
InstSp: E. Solheid
Observers: R. Pogge (Columbus); M. Wagner, J. Hill, O. Kuhn, M. Edwards, S. Walsh, A. Cardwell, M. de la Pena (Tucson)

Weather: Cirrus, especially at start and windy.

MODS2B: The fiber was connected to the PCI computer for remote testing from OSU. Elliott connected it to M2.BC. It must be reconnected to the PCI computer when we're done with MODS2 (we need to do some cals, still).

Start with MODS binocular observing.
  • 01:21 UT. Rick sent binocular acquisition on M33 target. Right preset failed because unable to find guide star. Pointing model not good enough at zenith (we were at ZD ~ 2.4 deg). SX ripped two times, cause unknown, but at that time of the night we had some strong winds: 5-7 m/s on secondary. Error message in log: Stop FLAO, Coils disabled. M33 is just too close to zenith and should be observed later. Moving on to the standard, for now.
  • G191B2B
    • Rick sent binocular acquisition. Observed - got only 2/3 of the MODS1 Red exposures as there was a comm glitch on this channel after the first exposure completed.
  • NGC1068
    • Acq images on blue/red saturated, modsAlign_v2 -B did not help, but adjusted scaling on modsDisp image to note X,Y peak and clicked "x" at this pixel position in modsAlign_v2.
    • Spectra are saturated around Halpha/[NII] on MODS2 - OK on MODS1.
  • M33 target
    • Used the 3x1200-sec script.
Michelle drives:

  • Rothberg Science MODS1 & 2
09:15 Preset. PCS problem (IT5817) which John solved. Resend the preset

09:25 Start Acquisition MODS1r [18] [19]; MODS2r [19] [20]. FWHM ~ 1.3"-1.7"

  • We see the bright star for alignment but we believe we also see the science target (given position of faint object at correct location). When we align on bright alignment star, target is NOT in slit MODS1r [20] and MODS2r [21] and it looks that alignment angle may actually be slightly off (10 pixels in X).
09:45 Undo offset and redo acquisition using same files, centering on science target. MODS1r [21] and MODS2r [22]
  • FWHM ~ 1.3"
    09:57 (1 X 1800s) science exposures started on MODS1 [r22; b10] and MODS2 [r23, b10]; FWHM 1.3-1.5". Emission lines detected.

    10:33 Acquisition for specphot mods1r [23]; mods2r [24]. FWHM ~ 1.3"-1.7"

  • G191B2B - collimation on MODS1 is slow and not converging. Active optics needed to be cleared on SX M1 and M2 simultaneously. Still the guide star looks poorly collimated, but this probe is close to the bottom edge of the field, where the WFS pupil is vignetted, and seeing is poor, exacerbating the problems collimating on the vignetted pupil image. Complete mods2r acquisition [25] while JHill tries to collimate mods1 by hand. Is successful
10:45 Start mods2 specphot 2 X 90s [r26-27; b11-12]

10:54 Redo mods1 acquisition for specphot [24-25] with improved collimation

10:56 Start mods1 specphot 2 X 90s [r26-27; b11-12]

Olga and JHill Drive:

  • Rothberg Science and LUCI1 & 2 Guiding check (JHill)
11:05 Reconfigure for LUCI1 and LUCI2 plus initial collimation and pointing check. LUCI1 pointing seems off, but we want to get to science so we\x92re going ahead with LUCI2

11:26 Preset LUCI2 science

11:30 GCS is having issues. See JHill e-mail.

11:45 Now some issues with LUCI2. Scheduler needs to be reset. Consulting with Dave on the mountain.

11:50 Further LUCI2 problems. Odd negative countdown on exposure time. Stopping script, reconfiguring for LUCI2 monocular for LBTB science.

12:00 Turnover to mountain.

21 January 2016: Calibrations for these MODS data.

  • MODS1 calibrations taken by LBTB observers on 20160117 (mods1x.20160116.*.fits) and by me on 20160117 (mods1x.20160117.*.fits and mods2x.20160117.*.fits), but the
MODS2 ones are all saturated and ruined by the sequencer problem.
  • comp lamps:
    • on 20160116: mods1r.2016.0116.0018,19,20.fits, mods1b.20160116.0012,13,14.fits and
    • on 20160117: mods1r.20160117.0001,2,3.fits and mods1b.20160117.0001,2,3.fits
  • pixflats: on 16: r 21-25 and b 15-19
    • then on 17: r 4-8 and b 4-8,9-13
  • 0.6" slit flats: r 26-31 and b 25-30
    • then on 17: r 9-13 and b 14-23;
  • 0.8" slit flats r 32-37 b31-36 and on 17, r 14-19 and b 24-33;
  • 1.0" slit flats r 38-43 b 37-46 and on 17, r 20-24 and b 34-43
  • biases: r 44-48

-- %USERSIG{OlgaKuhn - 2016-01-15}%

Comments

 
Topic revision: r4 - 23 Jan 2016, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback